OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x + 3 Is this the right answer y = 3x - 17 y = 3x - 5

OpenStudy (anonymous):

slope(m) = 3 (1,6) x = 1 and y = 6 y = mx + b (solving for b) 6 = 3(1) + b 6 = 3 + b 6 - 3 = b 3 = b your equation is : y = 3x + 3 --- you are correct :)